1) Calculate the difference in energy between the intrinsic Fermi level and the midgap energy at T = 300 K for Si.

2) Determine the electron and hole concentrations in Si at 300 K for NA = 1016 cm-3 and ND = 3 x 1015 cm-3 . You must derive the equations used to determine the concentrations


1
Expert's answer
2021-06-13T18:04:51-0400

1)

"n_i^2=N_cN_vexp[Eg\/KT]" from this We get, "n_i=1.5*10^{10}cm^{-3}."

At temperature T = 300K the values of effective density of states function in the conduction band ("N_c" ) and the effective density of states function in the valence band ("N_v" ) are "2.8*10^{25}m^{-3}" and "1.04*10^{25}m^{-3}" respectively. Assume the value of bandgap energy ("E_g" ) of silicon is 1.12 eV does not vary over this temperature range.

2)

The majority carrier electron concentration is

"n_o = \u00bd((10^{16} \u2013 3 * 10^{15}\n) + (((10^{16} \u2013 3 * 10^{15}\n)^2 + 4(1.5 * 10^{10})^2\n)^{1\/2}\n) \u2245 7 * 10^{15} cm^{-3}"

The minority carrier hole concentration is

"p_0 = \frac{ni_2}{n_0} = (1.5 * 10^{10}
)^2/(7 * 10^{15}) = 3.21 * 10^4 cm^{-3}"

If we assume complete ionization and "N_d - N_a >> n_i" , the majority carrier electron concentration is a very good approximation, just the difference between the donor and acceptor concentrations.